OpenStudy (anonymous):

p=2(l+w) solve for l

OpenStudy (anonymous):

p=2(l+w) p=2l+2w p-2w=2l (p-2w)/2 = l

OpenStudy (anonymous):

got it kriszer

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

that's over two right?

OpenStudy (anonymous):

yes kriszer a medal would be great

OpenStudy (anonymous):

Thanks